The table below shows the results of a survey in which researchers examined a child's IQ and the presence of a stupid gene is the child.

Gene Present Gene not Present Total
High IQ 33 11 44
Low IQ 12 27 39
Total 45 38 83

I.The Probability that a randomly selected child has Low IQ or Gene Present is closest to:

II. Given that the gene is present in a child, the probability that he/she will have Low IQ is closest to
Answer #1

From given table,

I

The Probability that a randomly selected child has Low IQ or Gene Present is ,

= P(low IQ) + P(Gene present) - P(low IQ and Gene present)

= (39 + 45 - 12) / 83

= 0.8675

II) P(low IQ | Gene present) = 12 / 45 = 0.2667